# Tie-Breaking Rule Based on Partial Proof of Work in a Blockchain

### v1

@@ V-A1 Theoretical Analysis
We theoretically demonstrate the effectiveness of our proposed method in suppressing Match against post-generated block. Specifically, we evaluate the proportion $\gamma$, of honest miners who mine on the attacker’s chain when the attacker performs Match against post-generated block. $\gamma$ is a general indicator. Most attacks that use intentional forks such as Selfish Mining also use intentional chain ties. Suppressing intentional chain ties is the same as suppressing $\gamma$.
@@
In a system in which all honest miners adopt our proposed method and an attacker initiates Match against post-generated block, the following conditions (Condition (1) and Condition (2)) can be considered as sufficient for the chain that is generated by honest miners to be selected among the chains in a tie.
@@
$=\exp(-p(\Sigma_{j\in V}P_{ij}M_{j}+(2\Delta_{B}+\Delta_{P})M_{all})$ (5)
$=\exp(-\Sigma_{j\in V}\frac{P_{ij}}{T}\frac{M_{j}}{M_{all}}+\frac{2\Delta_{B}+\Delta_{P}}{T})$ (6)
@@
From the above, the probability (Pr[Cond(1)]Pr[Cond(2)]) that an honest miner selects a block that is generated by an honest miner instead of one that is generated by the attacker satisfies the following inequality:
@@
The term $\Sigma_{j\in V}\frac{P_{ij}M_{j}}{M_{all}}$ in Equation 6 represents the hashrate-weighted average propagation time of a partial PoW when it is generated by miner $i$. It is widely known that the block propagation time and fork rate are deeply connected[18]. The relationship between the hashrate-weighted propagation time $T_{W}$ and the fork rate $F$ can be expressed as $F=T_{W}/T$[19]. The fork rate refers to the probability of another miner creating a block that forks the blockchain while a certain block is still propagating through the network. Therefore, knowing the fork rate enables the estimation of the hashrate-weighted propagation time of a block. Although determining the current value in Bitcoin is challenging, previous research[20] has shown the fork rate of Bitcoin to be approximately 0.41%. This implies a hashrate-weighted propagation time of approximately 2.4 s, which is significantly less than the measured 100th percentile propagation time of a block[16]. When applying our proposed method to Bitcoin, each partial PoW would be 80 bytes, which is much smaller than a block, suggesting that the hashrate-weighted propagation time of partial PoW could be even smaller. Consequently, the actual value of $Pr[Cond(2)]$ is likely to be closer to $\exp(\frac{2\Delta_{B}+\Delta_{P}}{T})$ than the value in inequality 7.
@@
In a system in which all miners adopt the proposed method, upon generating a block an attacker must publish shared partial PoWs that belong to the set corresponding to the shared partial PoW of the block. $2\Delta_{B}+2\Delta_{P}$ after the attacker publishes these partial PoWs, all honest miners include the attacker’s published partial PoWs in their shared partial PoW and engage in mining. Thus, if no new blocks are generated within $2\Delta_{B}+2\Delta_{P}$ after the attacker has generated a block, the honest miners will have at least one shared partial PoW of the same size as the attacker’s chain. In addition, a chain is randomly selected in the proposed method if the number of shared partial PoWs is equal. Considering these factors, the following inequality is established:
@@
Figures 2 and 3 depict the variations in $\gamma$ for different difficulty adjusters $n$, given the proportion of an attacker’s hashrate $\alpha$ of 0.5 and the average block generation interval $T$ of 600. Figure 2 shows the case in which $\Delta_{B}=\Delta_{P}=10$, whereas Figure 3 shows the results for $\Delta_{B}=\Delta_{P}=20$. As is evident from the figures, the value of $\gamma$ decreases as the difficulty adjuster $n$ increases. However, increasing $n$ also implies an increase in communication overhead. Thus, varying the difficulty adjuster according to the condition of the system is crucial. Additionally, the figures reveal that our proposed method reduces $\gamma$ more effectively than the random rule when $n$ is greater than 3. Specifically, for $n$ = 50 with $\Delta_{B}=\Delta_{P}=20$, $\gamma$ is reduced to 0.10118 or less, and for $\Delta_{B}=\Delta_{P}=20$, $\gamma$ is reduced to 0.15915 or less.
@@
Figure 4 depicts the effectiveness of the proposed method when the difficulty adjuster $n$ is fixed at 50 and the value $\frac{2\Delta_{B}+2\Delta_{P}}{T}$ is varied. As shown in the figure, the effectiveness of the proposed method decreases as $\frac{2\Delta_{B}+2\Delta_{P}}{T}$ increases. However, it is also apparent that the proposed method is more effective than the random approach for systems with shorter block generation intervals than Bitcoin, such as Litecoin (with a block generation interval of 2.5 min)[21]. The figure also indicates that the effectiveness of our method decreases as the proportion of the attacker’s hashrate increases. This diminished effectiveness is attributed to the reduction in the number of partial PoWs that are publicly available on the network. Increasing the difficulty adjuster $n$ can easily address this issue.
@@ V-A2 Simulation Analysis
We further evaluate our method through simulation analysis. The evaluation of the effectiveness of the method in Lemma 4 only considers situations in which all honest miners hold a shared partial PoW that is larger than the shared partial PoW of the attacker’s block. However, As explained after the proof of Lemma 4, a more stringent evaluation is possible by considering the hashrate-weighted propagation time of partial PoW, which is not considered in Lemma 4. In addition, the current assessment fails to consider the effect of each miner holding a shared partial PoW of the same size as the attacker’s block, as indicated in the evaluation presented in Lemma 5. Another overlooked aspect is the scenario in which honest miners successfully generate a block while some miners have not yet received the partial PoW that is published by the attacker. In such cases, miners, who haven’t received the attacker’s partial PoW, would select the block that is generated by an honest miner instead of the attacker’s block during a chain tie. This factor is also not considered in the evaluation presented in Lemma 5.
A theoretical evaluation that accounts for all factors is challenging. Therefore, we conducted a simulation experiment to assess the proposed method. The simulation allows us to consider both scenarios: when all honest miners possess a shared partial PoW that is larger than the attacker’s shared partial PoW, and when they have a shared partial PoW that is the same size as that of the attacker. We used SimBlock[22], which is a tool that can simulate a blockchain network. The default network parameters of SimBlock for the year 2019 were used. For simplicity, we did not apply Compact Block Relay[23], which is the default setting in SimBlock. The number of miners was set to 300, and the block generation interval $T$ was set to 600 s. The block size was adjusted to approximately 200,000 bytes and 500,000 bytes so that the 100%ile propagation time of the block was approximately 10 s and 20 s. For simplicity, the size of partial PoW was also set to 200,000 bytes and 500,000 bytes.
@@
The settings for the simulation described above were configured to be advantageous for the attacker. For example, the actual size of a partial PoW is 80 bytes, but we increased it to 200,000 and 500,000 bytes in our simulation. This enlargement inevitably lengthened the propagation time of the partial PoW, thereby reducing the likelihood that honest miners would hold a partial PoW larger than that of the attacker, compared to actual circumstances. Moreover, attackers in our simulation could instantly incorporate any received partial PoW into their shared partial PoW. This action could potentially lead to the inclusion of insufficiently shared partial PoWs in the shared partial PoW. However, in our setup, honest miners did not verify whether the partial PoWs were sufficiently shared. This increased the probability that the attacker’s shared partial PoW would be larger than that of the honest miners. In addition, although the attacker published their shared partial PoW upon block generation, this publication only occurred after a delay of $\Delta_{P}$. In reality, honest miners may generate a block during the propagation of shared partial PoW, and some honest miners may be unaware of this shared partial PoW. However, this risk was disregarded in our simulation.
Figures 2 and 3 show the results of the simulation experiments of 1000 times of chain ties. The data clearly show that the effectiveness of the proposed method, as measured by these simulations, exceeded the values obtained in the theoretical analysis. Specifically, when n = 50 and $\Delta_{B}=\Delta_{P}=10$, the value of $\gamma$ is 0.033. In contrast, for $\Delta_{B}=\Delta_{P}=20$, $\gamma$ increased to 0.07.
@@ V-B Suppression of Match against pre-generated block
We conducted an analysis with various settings for the unresponsive time and different values of $\Delta_{B}$ and $\Delta_{P}$. The results are shown in Figures 7 and 8. As seen in the figures, there was almost no difference in the relative revenue of the attacker between the SM and ESM. The required proportion of the attacker’s hashrate for successful SM was approximately $\alpha=0.32246$ with $s=20$ and $\Delta_{B}=\Delta_{P}=10$, and approximately $\alpha=0.31479$ with $s=40$ and $\Delta_{B}=\Delta_{P}=20$, whereas it was approximately $\alpha=0.32045$ and $\alpha=0.31055$, respectively, ESM. These results demonstrate that setting the acceptance window $w$ appropriately can mitigate Match against pre-generated block.
@@
VI Implementaion Details
VI-A Application to Existing Systems
In this section, we discuss the application of our proposed method to existing systems, particularly Bitcoin, and the associated effects.
The proposed method requires shared partial PoW to be embedded into blocks. Using the OP_RETURN[24] opcode is a known method for embedding shared partial PoW into a transaction. For example, an output can be reserved in the coinbase transaction specifically for embedding the shared partial PoW.
@@ VI-A Application to Existing Systems
Next, let’s consider the impact of our proposed method on the network bandwidth. Our method requires each miner to share a partial PoW. The way of sharing is arbitrary. For example, the same protocol used for propagating transactions can be used. When the proposed method is implemented in Bitcoin, a partial PoW is the block header, which is 80 bytes. This value is considerably smaller than the block size, and the increased bandwidth is acceptable. In addition, when a block is transmitted, it is necessary to send the set of partial PoWs corresponding to the shared partial PoW, which is approximately $80n$ bytes on average. The increased bandwidth can still be considered acceptable with an appropriate setting of $n$. Furthermore, the use of communication protocols that leverage set reconciliation, such as Graphene[26], can further reduce the bandwidth.
Additionally, our proposed method can be used as needed. The proposed method hardly changes the blockchain protocol, except for the tie-breaking rule. This allows for flexible responses such as using our proposed method only when the network is under attack and forks are frequent. For example, only when frequent forks are detected, each miner can send messages that require their neighbor nodes to send partial PoWs. This as-needed application improves communication traffic per miner.
@@ VI-D Clock Drift
Next, we reconsider the parameter settings for sufficiently shared partial PoW. If the time required for a valid partial PoW to be sufficiently shared is greater than $w+\Delta_{B}$, the partial PoW that is published by the attacker after an honest miner generates a block during a Match against post-generated block will not be regarded as sufficiently shared before the end of acceptance window of any honest miner. Therefore, similar to the parameter settings for the acceptance window, it is appropriate to consider a partial PoW as sufficiently shared after $2\Delta_{B}(1+D)$.
@@ VII Conclusion
We have proposed a novel last-genrated rule based on partial PoW for PoW blockchain systems. Our method is designed to suppress intentional chain ties that are caused by attackers. When applied to Bitcoin, our method increases the threshold of the proportion of the attacker’s hashrate that is necessary for selfish mining (SM) to $\alpha=0.31479$, regardless of the block propagation abilities of the attacker. Whereas existing methods require extensive updates or challenging assumptions for PoW blockchain systems, our method can be easily applied to existing systems including Bitcoin. In addition, we have proposed extended selfish mining (ESM) to exploit last-generated rules and demonstrated the importance of parameter settings in last-generated rules.

### v2

@@ V-A1 Theoretical Analysis
We theoretically demonstrate the effectiveness of our proposed method in suppressing Match against post-generated block. Specifically, we evaluate the proportion $\gamma$ of honest miners who mine on the attacker’s chain when the attacker performs Match against post-generated block. $\gamma$ is the most important general indicator. Most attacks that use intentional forks, such as Selfish Mining, also use intentional chain ties. Suppressing intentional chain ties is the same as suppressing $\gamma$.
@@
In a system in which all honest miners adopt our proposed method and an attacker initiates Match against post-generated block, the following conditions (Condition (1) and Condition (2)) can be considered sufficient for the chain that is generated by honest miners to be selected among the chains in a tie (Figure 3(b)).
@@
$=\exp(-p(\Sigma_{j\in V}P_{ij}M_{j}+(2\Delta_{B}+\Delta_{P})M_{all}))$ (5)
$=\exp(-\Sigma_{j\in V}\frac{P_{ij}}{T}\frac{M_{j}}{M_{all}}-\frac{2\Delta_{B}+\Delta_{P}}{T})$ (6)
@@
From the aforementioned, the probability (Pr[Cond(1)]Pr[Cond(2)]) that an honest miner selects a block that is generated by an honest miner instead of one that is generated by the attacker satisfies the following inequality:
@@
The term $\Sigma_{j\in V}\frac{P_{ij}M_{j}}{M_{all}}$ in Equation 6 represents the hashrate-weighted average propagation time of a partial PoW when it is generated by miner $i$. It is widely known that the block propagation time and fork rate are deeply connected[20]. The relationship between the hashrate-weighted propagation time $T_{W}$ and the fork rate $F$ can be expressed as $F=T_{W}/T$[21]. The fork rate refers to the probability of another miner creating a block that forks the blockchain while a certain block is still propagating through the network. Therefore, knowing the fork rate enables the estimation of the hashrate-weighted propagation time of a block. Although determining the current value of Bitcoin is challenging, previous research[22] has shown the fork rate of Bitcoin to be approximately 0.41%. This implies a hashrate-weighted propagation time of approximately 2.4 s, which is significantly less than the measured 100th percentile propagation time of a block[18]. When applying our proposed method to Bitcoin, each partial PoW would be 80 bytes, which is considerably smaller than a block, suggesting that the hashrate-weighted propagation time of partial PoW could be even smaller. Consequently, the actual value of $Pr[Cond(2)]$ is likely to be closer to $\exp(\frac{2\Delta_{B}+\Delta_{P}}{T})$ than the value in inequality 7.
@@
In a system in which all miners adopt the proposed method, upon generating a block an attacker must publish shared partial PoWs that belong to the set corresponding to the shared partial PoW of the block. $2\Delta_{B}+2\Delta_{P}$ after the attacker publishes these partial PoWs, all honest miners include the attacker’s published partial PoWs in their shared partial PoW and engage in mining. Thus, if no new blocks are generated within $2\Delta_{B}+2\Delta_{P}$ after the attacker has generated a block, the honest miners will have at least one shared partial PoW of the same size as the attacker’s chain (Figure 3(a)). In addition, a chain is randomly selected in the proposed method if the number of shared partial PoWs is equal. Considering these factors, the following inequality is established:
@@
Figures 4 and 5 depict the variations in $\gamma$ for different difficulty adjusters $n$, given the proportion of an attacker’s hashrate $\alpha$ of 0.5 and the average block generation interval $T$ of 600. Figure 4 shows the case in which $\Delta_{B}=\Delta_{P}=10$, whereas Figure 5 shows the results for $\Delta_{B}=\Delta_{P}=20$. As is evident from the figures, the value of $\gamma$ decreases as the difficulty adjuster $n$ increases. However, increasing $n$ also implies an increase in communication overhead. Thus, varying the difficulty adjuster according to the condition of the system is crucial. Additionally, the figures reveal that our proposed method reduces $\gamma$ more effectively than the random rule when $n$ is greater than 3. Specifically, for $n$ = 50 with $\Delta_{B}=\Delta_{P}=20$, $\gamma$ is reduced to 0.10118 or less, and for $\Delta_{B}=\Delta_{P}=20$, $\gamma$ is reduced to 0.15915 or less.
@@
Figure 6 depicts the effectiveness of the proposed method when the difficulty adjuster $n$ is fixed at 50 and the value $\frac{2\Delta_{B}+2\Delta_{P}}{T}$ is varied. As illustrated in the figure, the effectiveness of the proposed method decreases as $\frac{2\Delta_{B}+2\Delta_{P}}{T}$ increases. However, it is also apparent that the proposed method is more effective than the random approach for systems with shorter block generation intervals than Bitcoin, such as Litecoin (with a block generation interval of 2.5 min)[23]. The figure also indicates that the effectiveness of our method decreases as the proportion of the attacker’s hashrate increases. This diminished effectiveness is attributed to the reduction in the number of partial PoWs that are publicly available on the network. Increasing the difficulty adjuster $n$ can easily address this issue.
@@ V-A2 Simulation Analysis
We further evaluate our method through simulation analysis. The evaluation of the effectiveness of the method in Lemma 4 only considers situations in which all honest miners hold a shared partial PoW that is larger than the shared partial PoW of the attacker’s block. However, as explained after the proof of Lemma 4, a more stringent evaluation is possible by considering the hashrate-weighted propagation time of partial PoW, which is not considered in Lemma 4. In addition, the current assessment fails to consider the effect of each miner holding a shared partial PoW of the same size as the attacker’s block, as indicated in the evaluation presented in Lemma 5. Another overlooked aspect is the scenario in which honest miners successfully generate a block while some miners have not yet received the partial PoW that is published by the attacker. In such cases, miners, who have not received the attacker’s partial PoW, would select the block that is generated by an honest miner instead of the attacker’s block during a chain tie. This factor is also not considered in the evaluation presented in Lemma 5.
A theoretical evaluation that accounts for all factors is challenging. In addition, conducting experiments in a real network is also challenging. Therefore, we conducted a simulation experiment to assess the proposed method. The simulation allows us to consider both scenarios: when all honest miners possess a shared partial PoW that is larger than the attacker’s shared partial PoW, and when they have a shared partial PoW that is the same size as that of the attacker. We used SimBlock[24], which is a tool that can simulate a blockchain network. The default network parameters of SimBlock for the year 2019 were used. For simplicity, we did not apply Compact Block Relay[25], which is the default setting in SimBlock. The number of miners was set to 300, and the block generation interval $T$ was set to 600 s. The block size was adjusted to approximately 200,000 bytes and 500,000 bytes so that the 100%ile propagation time of the block was approximately 10 s and 20 s. For simplicity, the size of partial PoW was also set to 200,000 bytes and 500,000 bytes.
@@
The effectiveness of the proposed method in the real network is better than the simulation result from the viewpoint of suppressing intentional chain ties. This is because the settings for the simulation described above were configured to be advantageous for the attacker. For example, the actual size of a partial PoW is 80 bytes, but we increased it to 200,000 and 500,000 bytes in our simulation. This enlargement inevitably lengthened the propagation time of the partial PoW, thereby reducing the likelihood that honest miners would hold a partial PoW larger than that of the attacker, compared to actual circumstances. Moreover, attackers in our simulation could instantly incorporate any received partial PoW into their shared partial PoW. This action could potentially lead to the inclusion of insufficiently shared partial PoWs in the shared partial PoW. However, in our setup, honest miners did not verify whether the partial PoWs were sufficiently shared. This increased the probability that the attacker’s shared partial PoW would be larger than that of the honest miners. In addition, although the attacker published their shared partial PoW upon block generation, this publication only occurred after a delay of $\Delta_{P}$. In reality, honest miners may generate a block during the propagation of shared partial PoW, and some honest miners may be unaware of this shared partial PoW. However, this risk was disregarded in our simulation.
Figures 4 and 5 show the results of the simulation experiments of 1000 times of chain ties. The data clearly show that the effectiveness of the proposed method, as measured by these simulations, exceeded the values obtained in the theoretical analysis. Specifically, when n = 50 and $\Delta_{B}=\Delta_{P}=10$, the value of $\gamma$ is 0.033. In contrast, for $\Delta_{B}=\Delta_{P}=20$, $\gamma$ increased to 0.07.
@@ V-B Suppression of Match against pre-generated block
We analyzed with various settings for the unresponsive time and different values of $\Delta_{B}$ and $\Delta_{P}$. The results are shown in Figures 9 and 10. As seen in the figures, there was almost no difference in the relative revenue of the attacker between the SM and ESM. The required proportion of the attacker’s hashrate for successful SM was approximately $\alpha=0.32246$ with $s=20$ and $\Delta_{B}=\Delta_{P}=10$, and approximately $\alpha=0.31479$ with $s=40$ and $\Delta_{B}=\Delta_{P}=20$, whereas it was approximately $\alpha=0.32045$ and $\alpha=0.31055$, respectively, ESM. These results demonstrate that setting the acceptance window $w$ appropriately can mitigate Match against pre-generated block.
@@
VI Implementation Details
VI-A Application to Bitcoin
In this section, we discuss the application of our proposed method to existing systems, particularly Bitcoin, and the associated effects. The method that we discuss here is the n-variable method.
Figure 12: Data structure of a block with the proposed method. The data enclosed in a red frame represent new data required for the specifications of the proposed method. Each partial PoW is the block header whose hash value meets the difficulty determined by the difficulty adjuster.
Figure 13: Flow of sharing partial PoWs. When a miner has detected frequent forks, it sends the request message to send partial PoWs to its neighbor nodes. Miners do not have to send the block corresponding to a partial PoW. This is because miners can verify that the partial PoW has valid proof of work only with itself.
Embedding shared partial PoWs and mining. The proposed method requires shared partial PoW to be embedded into blocks (Figure 12). Using the OP_RETURN[26] opcode is a known method for embedding shared partial PoW into a transaction. For example, an output can be reserved in the Coinbase transaction specifically for embedding the shared partial PoW.
@@ VI-A Application to Bitcoin
Committing to the difficulty adjuster. The n-variable method method allows miners to choose the difficulty adjuster $n$. This stabilizes the supply of partial PoW regardless of the number of miners who use the method.
Each miner sets the difficulty adjuster $n$ in response to the amount of partial PoW currently published during mining. In other words, each miner commits to the difficulty adjuster $n$. The way of commitment is very simple. We use the timestamp of a block header. This timestamp is Unix time and proceeds per second. The determination of $n$ relies on interpreting the initial bits of this timestamp (Figure 12). For example, if the first 2 bits are 00, $n=50$. If the first 2 bits are 01, $n=100$. If the first 2 bits are 10, $n=150$. If the first 2 bits are 11, $n=200$.
Each miner sets $n=50$ when the supply of partial PoW is sufficient, and vice versa. We consider the supply of partial PoW to be sufficient when there are more than 25 partial PoWs per block on average, which corresponds to a case where the proportion of hashrate of the attacker is $0.5$ and all honest miners adopt the proposed method and set $n=50$. For example, each miner first sets $n=200$. If there are enough partial PoWs, say more than 100 partial PoWs per block on average, it resets $n=150$. If there are 20 partial PoWs per block on average, it resets $n=200$ again.
Miners compare the sum of the reciprocals of difficulty adjusters corresponding to each partial PoW in a chain tie with the n-variable method while they compare the number of shared partial PoW with the n-fixed method. If miners cannot find shared partial PoW in the block, they assign a value of 0 to the block.
The impact of the variable difficulty adjuster is negligible. This is because it only affects the initial bits of the timestamp and a few seconds. Miners cannot distinguish the effect of the method from that of the block propagation time.
In addition to that, it is obvious that this n-variable method method has both forward compatibility and backward compatibility. Miners who do not adopt the method cannot distinguish blocks that are generated by themselves from blocks that are generated by miners who adopt the method. Similarly, miners who adopt the method cannot distinguish blocks that are generated by themselves from blocks that are generated by miners who do not adopt the method.
Sharing partial PoWs. Our method requires each miner to share a partial PoW. The way of sharing is arbitrary. For example, the same protocol used for propagating transactions can be used.
We can share partial PoWs as needed (Figure 13). The proposed method hardly changes the blockchain protocol, except for the tie-breaking rule. This allows for flexible responses such as using our proposed method only when the network is under attack and forks are frequent. For example, only when frequent forks are detected, each miner can send messages that require their neighbor nodes to send partial PoWs. This as-needed application improves communication traffic per miner.
Next, let’s consider the impact of our proposed method on the network bandwidth. When the proposed method is implemented in Bitcoin, a partial PoW is the block header, which is 80 bytes. That means the expected communication traffic is approximately $80n$ bytes per miner over 10 min. Considering that each miner shares a block’s worth of transactions every 10 minutes, the sharing cost of partial PoW is small. This value is considerably smaller than the block size, and the increased bandwidth is acceptable. In addition, when a block is transmitted, it is necessary to send the set of partial PoWs corresponding to the shared partial PoW, which is approximately $80n$ bytes on average. The increased bandwidth can still be considered acceptable with an appropriate setting of $n$. Furthermore, the use of communication protocols that leverage set reconciliation, such as Graphene[28], can further reduce bandwidth consumption.
@@ VI-B Partial Application
The proposed method is a tie-breaking rule. Hence, miners who adopt the proposed method do not generate stale blocks resulting from partial application. This property does not hold in the case of the application of fork choice rules such as GHOST [17] and weighted FRP[6].
@@ VI-D Clock Drift
Next, we reconsider the parameter settings for sufficiently shared partial PoW. If the time required for a valid partial PoW to be sufficiently shared is greater than $w+\Delta_{B}$, the partial PoW that is published by the attacker after an honest miner generates a block during a Match against post-generated block will not be regarded as sufficiently shared before the end of the acceptance window of any honest miner. Therefore, similar to the parameter settings for the acceptance window, it is appropriate to consider a partial PoW as sufficiently shared after $2\Delta_{B}(1+D)$.
@@ VII Conclusion
We have proposed a novel last-generated rule based on partial PoW for PoW blockchain systems. Our method is designed to suppress intentional chain ties that are caused by attackers. When applied to Bitcoin, our method increases the threshold of the proportion of the attacker’s hashrate that is necessary for selfish mining (SM) to $\alpha=0.31479$, regardless of the block propagation abilities of the attacker. Whereas existing methods require extensive updates or challenging assumptions for PoW blockchain systems, our method can be easily applied to existing systems including Bitcoin. In addition, we have proposed extended selfish mining (ESM) to exploit last-generated rules and demonstrated the importance of parameter settings in last-generated rules.